# Exploring Holomorphic Retracts

### v6

@@ Abstract.
The purpose of this article is towards systematically characterizing holomorphic retracts of domains of holomorphy; to begin with, bounded balanced pseudoconvex domains $B\subset\mathbb{C}^{N}$. Specifically, we show that every (holomorphic) retract of $B$ passing through its center (i.e., the origin), is the graph of a holomorphic map over a linear subspace of $B$. If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics; the necessity of the boundary points to be extreme points is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well. This revisit of Vesentini’s result leads to a small modification that is useful to provide interesting examples; specifically, in completely determining the retracts through the origin of certain balanced domains with analytic varieties in their boundaries. The converse of Vesentini’s theorem is established and general results characterizing the non-linear retracts through the origin are attained (for fairly general domains in this context) when the boundary fails to have extreme points. As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ which is submersive at its center, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$. To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of strictly convex balls, with respect to various norms and obtain a complete description of retracts passing through the origin, generalizing that of the polydisc due to Heath and Suffridge and giving a simpler proof as well. The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$. Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case. Indeed, for the simplest possible case here namely the ‘$\ell^{q}$-ball’ for $10<q<$ 1, we obtain a complete list of all retracts through the origin of each possible dimension; this enables an illustration of applying retracts to establishing biholomorphic inequivalences. To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof. Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial. We conclude with some expositions about retracts of $\mathbb{C}^{2}$.
@@
Union problem, Kobayashi hyperbolic, Kobayashi corank one, balanced pseudoconvex domains
@@ 1. Introduction
We begin with remarks for motivation and background, possibly a bit elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to. This will also simultaneously help us set up notations and fix terminologies. We emphasize that this article is a survey in parts, striving to give references for the known results (to the best of our knowledge). By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map. More precisely, we shall be interested in holomorphic retracts i.e., the image of a holomorphic retraction which by definition, is a holomorphic self-map $\rho$ of a domain $D\subset\mathbb{C}^{N}$ (or more generally, a complex manifold) satisfying the idempotent condition $\rho^{\circ 2}:=\rho\circ\rho=\rho$. This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points of $D$ fixed by $\rho$. A pair of extreme or trivial examples of retractions (resp. retracts) of any complex manifold are the identity map and constant maps (resp. $D$ and its points); obviously, we would not like to talk of them any further and thereby shall be referred to as trivial retracts henceforth. For instance, with this terminology and understanding, we shall say that in dimension $N=1$, there are no non-trivial retracts. This means that for any domain $D$ in $\mathbb{C}$ or any Riemann surface, any holomorphic self-map $\rho$ satisfying $\rho^{\circ 2}=\rho$ is either a constant map or else, the identity map on $D$. Indeed, this follows from the open mapping theorem together with, the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$. In other words, holomorphic retracts are interesting only when the dimension $N$ gets bigger than one. Now in multidimensional complex analysis, a lot of effort has gone into the study of biholomorphic and more generally, proper (holomorphic [^1] Unless otherwise mentioned, all maps in this article are to be understood to be holomorphic.) maps, as is natural. To mention just one or two of the first basic facts about proper maps between a pair of domains in $\mathbb{C}^{N}$, they are known to be always surjective and for many a class of domains $D$ with ‘nice’ boundaries, proper self-maps turn out to be automorphisms of $D$. Though absolutely trivial to check, let us mention for contrast with the just mentioned facts about proper maps, that: non-trivial retractions are never surjective (and never are they, automorphisms!); as one another sharp contrast, proper maps never decrease dimension, whereas non-trivial retractions always decrease dimension. In other words, retractions could be viewed as being, loosely speaking, at the ‘opposite end of the spectrum’ when compared to proper maps.
@@
While much work has been carried out on proper holomorphic maps and their images, far less investigation has been done about retracts. Infact, in addition to (the foregoing remarks and) their role of in certain fundamental mapping problems above, retracts are of significance in complex dynamics as well. To just be specific by providing a standard result in this regard, we mention the following theorem which shows how retracts arise in the study of iterates of a holomorphic self-map of taut manifolds. Needless to say, this theorem has been applied to the study of the special case of proper holomorphic self-maps. In light of the the stark contrasts between proper maps and retraction maps mentioned above, the connection amongst these maps, is thereby all the more interesting.
@@
[…]
@@
Let us also mention here for clarity that our focus will be to determine the form of the retracts $Z=F(X)$ rather than retractions mappings $F$. This owes to the fact that for a given retract $Z$, there are in general plenty of retractions all of whose images are the same $Z$. This is manifestly brought out already in the simplest of cases for $X$ namely, when $X$ is the bidisc $\Delta^{2}$, a description of whose retracts was attained (indeed for $\Delta^{N}$) in 1981 by Heath and Suffridge in [Heath_Suff]. The case of the Euclidean ball dealt with in 1974 by Suffridge in [Suffridge-Ball] is exposited well in Rudin’s book [Rud_book_Fn_theory_unitball] (we shall not detail the timeline of the results, beyond the most basic cases of the ball and polydisc). As highlighted in the introduction of this article by Heath – Suffridge, there is a wide range of holomorphic retractions of the bidisc with the same image, which we recall here for some later significance for our article. Let $10<t<$ 1 and pick an arbitrary holomorphic function $f:\Delta^{2}\to\mathbb{C}$ which is bounded above in modulus by $t(1-t)/2$. Then the mapping on $\Delta^{2}$, defined by
@@
for any fixed $\theta\in\mathbb{R}$, is a holomorphic retraction of $\Delta^{2}$ onto $\left\{z\left(1,e^{-i\theta}\right):|z|<1\right\}$, as is checked by a straightforward computation which we skip here, to get ahead with building a perspective in this introductory section. Note that though such a retraction may well be highly non-linear and have a complicated expression as just exemplified, its image (regardless of what $f$ is), is just a linear copy of the disc. This finishes the demonstration that for a given retract $Z$, there may very well be plenty of (as remarked in [Heath_Suff] ‘quite complicated’!) retractions all of whose images are the same $Z$. At this juncture, it may tempting to ask as to whether this phenomenon of the existence of rather complicated holomorphic retractions with the same image, owes to the polydisc being a domain that reduces to a product of one-dimensional discs, thereby in-particular, giving rise to a plethora of analytic discs in the boundary. That this is not the case is borne out by the fact that the same is true (in-fact, in some sense all the more! – see below) for Euclidean balls, as is recorded in chapter-8 of Rudin’s book [Rud_book_Fn_theory_unitball] wherein complicated retraction maps of the Euclidean ball $\mathbb{B}^{2}\subset\mathbb{C}^{2}$, all with the same image, are exemplified. Indeed to recall this briefly here, pick any $f\in H^{\infty}(\mathbb{B}^{2})$ of the form
@@
where the $g_{j}$’s are arbitrary elements of $H^{\infty}(\mathbb{B}^{2})$; then the associated self-map $F$ of $\mathbb{B}^{2}$ given by $F(z,w)=\left(f(z,w),0\right)$ is a retraction map, as simply follows from the fact that $f(z,0)\equiv z$. Its image is $\pi_{1}(\mathbb{B}^{2})=\Delta$, where $\pi_{1}$ is the linear projection onto the first factor. Hence again, while $f$ and thereby $F$ may well be be very complicated, the retract here is just a disc. We are therefore motivated to focus on retracts first more than retraction mappings.
@@
Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; but the converse need not be true in general. But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map of any taut manifold $X$, is contained within a non-trivial retract. Furthermore, there are results for an exact equality rather than just containment, if so sought for: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain. Whether or not fixed point sets of a holomorphic self map is the retract of possibly some other holomorphic self map of $X$, a definitive basic feature that all retracts have in common is that every retract $Z$ has the extension property in the sense that every holomorphic function on $Z$ admits (just by composition with the retraction mapping $X$ onto $Z$) a holomorphic extension to $X$. When the given function on $Z$ is bounded, we also have (by the same means of composition with the retraction map), a holomorphic extension on $X$ with the same sup-norm and thus retracts form the simplest class of subvarieties of $X$ with the norm-preserving extension property to $X$. Whether retracts are the only subvarieties with this property was examined and shown to be so, in many special cases by Kosiński and McCarthy in the relatively recent work [Kosnski] wherein counterexamples have also been given. All this is only towards indicating the significance and motivation for the study of holomorphic retracts, in general. Our first goal for this article is rather specific and it is: to obtain a description of the holomorphic retracts of pseudoconvex bounded balanced domains passing through its center, namely the origin in $\mathbb{C}^{N}$. This itself is noteworthy; if not for anything else, this suffices for an important application to attain (employing the above theorem 1.1), the fairly recent result of [JJ] together with [Pranav_JJ] that: every proper holomorphic self-map of any smoothly bounded balanced pseudoconvex domain of finite type is an automorphism. We shall not however elaborate upon this here.
@@
As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E. Vesentini already proved the following fundamental result about the form of the retracts of certain convex domains. More important than the various hypotheses of this result, as pointed out by J. P. Vigue in [Vigue_fxdpnt_cnvx], Vesentini’s approach makes ‘heavy’ use of complex geodesics. We shall to begin with, give a relatively elementary proof of this result without using complex geodesics. It may be noted that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal; thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$, but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal but indeed $\mathbb{C}$-extremal. The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries. The necessity of such extremality hypotheses is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well.
@@
Here and throughout the article, by a linear subspace $D_{L}$ of a balanced domain $D\subset\mathbb{C}^{N}$, we mean that $D_{L}$ is of the form $D_{L}=D\cap L$, where $L$ is a linear subspace of $\mathbb{C}^{N}$. Likewise, by an affine linear subspace $A^{\prime}$ of $D$, we mean $A^{\prime}=A\cap D$ where $A$ is an affine linear subspace of $\mathbb{C}^{N}$. To recall a relevant fact in this parlance, every retract of a ball (centered at the origin, as always in this article) linearly equivalent to the standard Euclidean ball $\mathbb{B}^{N}$ is affine linear. We postpone rendering clarity about a potential confusion about the notion of linear retracts to section 3; likewise, for definitions of a couple of terms as well, in order to focus more in this introductory section, on the narrative of our results and work about retracts in perspective of what is already known.
Now, the condition of convexity on $D$, which entails a certain condition to be satisfied by the entire boundary $\partial D$, can be dropped in various ways. To begin with here, let us only remark that we can do so in the following manner; this will actually be useful in dealing with certain concrete examples in the sequel (example 8.7 in section 8). This is one another major reason for in revisiting this result of Vesentini. Specifically, we shall show the following without using complex geodesics as alluded to earlier.
Theorem 1.3.
(A small modification of Vesentini’s result): Let $D$ be any bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$. Suppose $Z$ is a retract through the origin with the property that: $D\cap T_{0}Z$ is convex with points of $\partial D\cap T_{0}Z$ being $\mathbb{C}$-extremal boundary points of $D$; then $Z$ is a linear subspace of $D$. In particular, every $1$-dimensional retract through the origin of a bounded balanced pseudoconvex domain $D$ with $\mathbb{C}$-extremal boundary, is linear.
Our inclusion of this result first towards a relatively self-contained exposition, has other more important motivations than just a more elementary proof: to furnish examples of balanced non-convex (but pseudoconvex) domains possessing non-trivial analytic varieties in the boundary, wherein the retracts through the center can be completely determined. This is accomplished in section 8 wherein the above modified version of Vesentini’s result has also a role to play in other propositions there. Also in general, we would not like to be amiss to observe associated facts which are not explicitly recorded in the literature but can be attained by reading proofs of results carefully – these are mostly exposited in the ‘preliminaries’ section; more importantly, such observations are actually useful for us much later in the sequel. While we relegate such observations to later, one that we think must be mentioned right away here, is the affirmative answer to the question about the very existence of atleast one retract through the origin of an arbitrary bounded balanced domain of holomorphy without any further assumptions. This is verified in proposition 2.35. Getting back to the above theorem let us now mention for instance, the apparently strong condition (when $\dim(Z)\geq 2$) on the convexity of $D\cap T_{0}Z$ in the above theorem can be dropped while attaining the same conclusion; specifically, we have the following.

### v7

@@ Abstract.
The purpose of this article is towards systematically characterizing holomorphic retracts of domains of holomorphy; to begin with, bounded balanced pseudoconvex domains $B\subset\mathbb{C}^{N}$. Specifically, we show that every (holomorphic) retract of $B$ passing through its center (i.e., the origin), is the graph of a holomorphic map over a linear subspace of $B$. If $B$ is further assumed to be convex with $\mathbb{C}$-extremal boundary, then every such retract is linear – this is a well-known result due to Vesentini but we give here an elementary proof without making use of complex geodesics which works when convexity is dropped and $\mathbb{C}$-extremality is replaced by holomorphic extremality. The converse of Vesentini’s theorem is also established and general results characterizing the non-linear retracts through the origin are attained (for fairly general domains in this context) when the boundary fails to have extreme points. As for retracts not passing through the origin, we obtain the following result: if $B$ is a strictly convex ball with respect to some norm on $\mathbb{C}^{N}$, and $\rho$ any holomorphic retraction map on $B$ whose derivative at the center is a norm-one projection, then $Z=\rho(B)$ is the graph of a holomorphic map over a linear subspace of $B$. To deal with a case when $\partial B$ may fail to have sufficiently many extreme points, we consider products of bounded balanced domains of holomorphy with holomorphic extreme boundaries and obtain a complete description of retracts passing through the origin. The conditions of this result can indeed be verified to be applied to solve a special case of the union problem with a degeneracy, namely: to characterize those Kobayashi corank one complex manifolds $M$ which can be expressed as an increasing union of submanifolds which are biholomorphic to a prescribed homogeneous bounded balanced domain in $\mathbb{C}^{N}$ (a description of all of its retracts being attained by the aforementioned result). Along the way, we discuss examples of non-convex but pseudoconvex bounded balanced domains in which most one-dimensional linear subspaces fail to be holomorphic retracts, in good contrast to the convex case. Indeed, for the simplest possible case here namely the $\ell^{q}$-ball’ for $0<q<1$ as well as its ‘anisotropic’ analogues, we obtain a complete list of all retracts through the origin of each possible dimension; whether $q$ is less than $1$ or not, retracts through the origin are shown to be biholomorphic copies of lower dimensional $\ell^{q}$-balls. The same is done also for a class of domains lying at the other extreme in their Levi geometry but nevertheless share with $\ell^{q}$-balls the feature of their boundaries being devoid of complex line segments; namely, balanced $\mathbb{C}$-extremal analytic polyhedra. In the course of achieving this, we prove a generalization of Mazet’s Schwarz lemma for pseudoconvex balanced domains. The above mentioned determination of retracts enables an illustration of applying retracts to establishing biholomorphic inequivalences. To go beyond balanced domains, we then first obtain a complete characterization of retracts of the Hartogs triangle and ‘analytic complements’ thereof. Thereafter, we drop the assumption of boundedness as well, to attain similar characterization results for domains which are neither bounded nor topologically trivial. We conclude with remarks about retracts of $\mathbb{C}^{2}$ including an alternative proof about the characterization of its polynomial retracts.
@@
Holomorphic Retracts, Balanced pseudoconvex domains, Schwarz lemma, Biholomorphic inequivalences, Union Problem, Kobayashi hyperbolic, Caratheodory geodesics
@@ 1. Introduction
We begin with remarks for motivation and background, possibly somewhat elaborate; this owes to the unavailability of a survey or expository article on holomorphic retracts to refer the reader to. This will also simultaneously help us set up notations and fix terminologies. While the main purpose of this article is to contribute original results, we wish to put our results systematically in connection with what is known for reasons already indicated. In this regard, we emphasize that this article is a survey in parts as well, striving to give references for the known results (to the best of our knowledge). By a retraction we shall mean an idempotent map, whereas by a retract, we mean the image of such a map. More precisely, we shall be interested in holomorphic retracts i.e., the image of a holomorphic retraction which by definition, is a holomorphic self-map $\rho$ of a domain $D\subset\mathbb{C}^{N}$ (or more generally, a complex manifold) satisfying the idempotent condition $\rho^{\circ 2}:=\rho\circ\rho=\rho$. This means that $\rho$ acts as the identity map on $Z=\rho(D)$ which shall be our notation for a retract throughout this article; a trivia worth noting is that $Z$ is precisely the set of points in $D$ fixed by $\rho$. A pair of extreme or trivial examples of retractions (resp. retracts) of any complex manifold are the identity map and constant maps (resp. $D$ and its points); obviously, we would not like to talk of them any further and thereby shall be referred to as trivial retracts henceforth. For instance, with this terminology and understanding, we shall say that in dimension $N=1$, there are no non-trivial retracts. This means that for any domain $D$ in $\mathbb{C}$ or any Riemann surface, any holomorphic self-map $\rho$ satisfying $\rho^{\circ 2}=\rho$ is either a constant map or else, the identity map on $D$. Indeed, this follows from the open mapping theorem together with the idempotency condition which implies that $\rho(D)$ is a closed subset of $D$. In other words, holomorphic retracts are interesting only when the dimension $N$ gets bigger than one. Now in multidimensional complex analysis, a lot of effort has gone into the study of biholomorphic and more generally, proper (holomorphic [^1] Unless otherwise mentioned, all maps in this article are to be understood to be holomorphic.) maps, as is natural. To mention just one or two of the first basic facts about proper maps between a pair of domains in $\mathbb{C}^{N}$, they are known to be always surjective and for many a class of domains $D$ with ‘nice’ boundaries, proper self-maps turn out to be automorphisms of $D$. Though absolutely trivial to check, let us mention for contrast with the just mentioned facts about proper maps, that: non-trivial retractions are never surjective (and never are they, automorphisms!); as one another sharp contrast, proper maps never decrease dimension, whereas non-trivial retractions always decrease dimension. In other words, retractions could be viewed as being, loosely speaking, at the ‘opposite end of the spectrum’ when compared to proper maps.
@@
While much work has been carried out on proper holomorphic maps and their images, far less investigation has been done about retracts. Infact, in addition to (the foregoing remarks and) their role in certain fundamental mapping problems indicated above, retracts are of significance in complex dynamics as well. To just be specific by providing a standard result in this regard, we mention the following theorem which shows how retracts arise in the study of iterates of a holomorphic self-map of taut manifolds. Needless to say, this theorem has been applied to the study of the special case of proper holomorphic self-maps. In light of the the stark contrasts between proper maps and retraction maps mentioned above, the connection amongst these maps, is thereby all the more interesting.
@@
The manifold $M$ occurring in the above theorem – referred to as the limit manifold – is, as is evident, a holomorphic retract of $X$. Determining all holomorphic retracts of taut manifolds in general is way too difficult. We shall therefore be taking up first, the question of describing the holomorphic retracts in what may seem to be a rather small class, namely that of bounded balanced pseudoconvex domains $D$ in $\mathbb{C}^{N}$; not all such domains need be taut, but all bounded balanced domains which are taut, are contained within this class (this owes to results of Barth [Barth]). Whether or not $D$ is taut, a first general fact to be noted is that every retract of $D$ is a Stein manifold. This follows from the basic fact that retracts of (connected) complex manifolds are closed connected complex submanifolds and if the manifold is a Stein manifold, then so is every of its retracts; as our class of domains of interest is contained within that comprising domains of holomorphy in $\mathbb{C}^{N}$ which are the simplest instances of Stein manifolds, the retracts in this article, all have this property. Perhaps before the just mentioned inheritance of the property of being a Stein manifold, we must recall in a sentence here the elementary fact that various topological properties such as simply connectedness or more generally the triviality of any of the homotopy or homology groups (of the ambient space) is inherited under passage to a retract. Indeed in general, if $R\xhookrightarrow{\iota}X$ denotes a (continuous) retraction map on a topological space $X$, then the homomorphism induced by the inclusion $\iota$ at the level of each homotopy or homology group, is injective since $r_{*}\circ\iota_{*}\equiv{\rm identity}$, in standard notations; this may be paraphrased as the general observation that topological complexity cannot increase for a retract (and our results will establish certain finer analogues of this in several special cases in the complex analytic category). While this means that all such groups must be trivial for every retract of a contractible space $X$, we cannot deduce as readily out of this, the fact that contractibility is inherited as well. Indeed, this inheritance follows directly from the definition of contractibility and the only reason we are spending words here to spell this out is because we shall be dealing with balanced domains for the most part, and all such domains are evidently contractible. It may be worth remarking once here that as soon as we restrict our considerations to the category of smooth manifolds, we gain easily defined properties that can be used to distinguish a pair of retracts, namely for instance, the dimension of a retract i.e., as soon as we know that a given pair of retracts of a contractible manifold are of different dimensions, we are ascertained that they are diffeomorphically inequivalent. This is just to indicate that classifying retracts of a manifold which is topologically trivial (here, contractible) is not an intractable or meaningless exercise. We shall in-fact be focussing our attention in good measure on bounded balanced domains of holomorphy. Let us only remark here as to why we do not replace boundedness by the apparently more general assumption of Kobayashi hyperbolicity: a balanced domain is hyperbolic if and only if it is bounded ([Kodama]) whereby in particular, the class of hyperbolic balanced convex domains (in $\mathbb{C}^{N}$) is precisely the class of bounded balanced convex domains. Now, boundedness implies hyperbolicity and it is known that a convex domain is taut if and only if it is hyperbolic  [^2] In such a case the convex domain is biholomorphic to a bounded domain but not necessarily to a bounded convex domain, as shown by Zimmer [Zim]. ([Bracci-Saracco])). It follows that the foregoing theorem in conjunction with our results would be useful towards obtaining a description of the limit manifolds of (proper) holomorphic self-maps of those bounded balanced convex domains that we consider in this article, in particular. Even within this class this question of determining all retracts has indeed been of interest but has been rather a difficult one and a complete description of all retracts is available only for the Euclidean ball $\mathbb{B}$, the polydisc $\Delta^{N}$ and perhaps their Cartesian products (herein and henceforth, all balls are centered at the origin and of unit radius with respect to some norm on $\mathbb{C}^{N}$; when the norm is the $l^{2}$ norm, the corresponding ball referred to as the Euclidean ball is denoted $\mathbb{B}^{N}$ or $\mathbb{B}$ for short; $\Delta$ denotes the standard unit disc in $\mathbb{C}$).
@@
Let us also mention here for clarity that our focus will be to determine the form of the retracts $Z=F(X)$ rather than retractions mappings $F$. This owes to the fact that for a given retract $Z$, there are in general plenty of retractions all of whose images are the same $Z$. This is manifestly brought out already in the simplest of cases for $X$ namely, when $X$ is the bidisc $\Delta^{2}$, a description of whose retracts was attained (indeed for $\Delta^{N}$) in 1981 by Heath and Suffridge in [Heath_Suff]. The case of the Euclidean ball dealt with in 1974 by Suffridge in [Suffridge-Ball] is exposited well in Rudin’s book [Rud_book_Fn_theory_unitball]; we shall not detail the timeline of the results, beyond the most basic cases of the ball and polydisc. As highlighted in the introduction of this article by Heath – Suffridge, there is a wide range of holomorphic retractions of the bidisc with the same image, which we recall here for some later significance for our article. Let $0<t<1$ and pick an arbitrary holomorphic function $f:\Delta^{2}\to\mathbb{C}$ which is bounded above in modulus by t(1t)/2t(1-t)/2. Then the mapping on $\Delta^{2}$, defined by
@@
for any fixed $\theta\in\mathbb{R}$, is a holomorphic retraction of $\Delta^{2}$ onto $\left\{z\left(1,e^{-i\theta}\right):|z|<1\right\}$, as is checked by a straightforward computation which we skip here, to get ahead with building a perspective in this introductory section. Note that though such a retraction may well be highly non-linear and have a complicated expression as just exemplified, its image (regardless of what $f$ is), is just a linear copy of the disc. This finishes the demonstration that for a given retract $Z$, there may very well be plenty of (as remarked in [Heath_Suff] ‘quite complicated’!) retractions all of whose images are the same $Z$. At this juncture, it may be tempting to ask as to whether this phenomenon of the existence of rather complicated holomorphic retractions with the same image, owes to the polydisc being a domain that reduces to a product of one-dimensional discs, thereby in-particular, giving rise to a plethora of analytic discs in the boundary. That this is not the case is borne out by the fact that the same is true (in-fact, in some sense all the more! – see below) for Euclidean balls, as is recorded in chapter-8 of Rudin’s book [Rud_book_Fn_theory_unitball] wherein complicated retraction maps of the Euclidean ball $\mathbb{B}^{2}\subset\mathbb{C}^{2}$, all with the same image, are exemplified. Indeed to recall this briefly here, pick any $f\in H^{\infty}(\mathbb{B}^{2})$ of the form
@@
where the $g_{j}$’s are arbitrary elements of $H^{\infty}(\mathbb{B}^{2})$; then the associated self-map $F$ of $\mathbb{B}^{2}$ given by $F(z,w)=\left(f(z,w),0\right)$ is a retraction map, as simply follows from the fact that $f(z,0)\equiv z$. Its image is $\pi_{1}(\mathbb{B}^{2})=\Delta$, where $\pi_{1}$ is the standard linear projection onto the first factor. Hence again, while $f$ and thereby $F$ may well be be very complicated, the retract here is just a disc. We are therefore motivated to focus on retracts, first more than retraction mappings.
@@
Let us then begin with the obvious observation that retracts are always fixed point sets of some holomorphic map; the converse however is not true in general. But then again, an easy but noteworthy consequence of theorem 1.1 is that the fixed point set of every holomorphic self-map on any taut manifold $X$, is contained within a non-trivial retract333these little claims will be justified in the preliminaries and subsequent sections.. Furthermore, there are results for an exact equality rather than just containment: as shown in [Mazet_Vigue_fxdpnt], [Thai_Fxdpnt_cnvx] among others, the fixed point set of any holomorphic self map of any hyperbolic convex domain in $\mathbb{C}^{N}$ – whenever it is non-empty – is also a retract of the domain. Whether or not fixed point sets of a holomorphic self map is the retract of possibly some other holomorphic self map of $X$, a definitive basic feature that all retracts have in common is that every retract $Z$ has the extension property in the sense that every holomorphic function on $Z$ admits (just by composition with the retraction mapping $X$ onto $Z$) a holomorphic extension to $X$. When the given function on $Z$ is bounded, we also have (by the same means of composition with the retraction map), a holomorphic extension on $X$ with the same sup-norm and thus retracts form the simplest class of subvarieties of $X$ with the norm-preserving extension property to $X$. Whether retracts are the only subvarieties with this property was examined and shown to be so, in many special cases by Kosiński and McCarthy in the relatively recent work [Kosnski] wherein counterexamples have also been given. This is the only sentence that we spend in recalling the recent history about the works on retracts in general, all other citations being actually used in one way or the other in the sequel. All this is only towards indicating the significance and motivation for the study of holomorphic retracts, in general. Our first goal for this article is rather specific and it is: to obtain a description of the holomorphic retracts of pseudoconvex bounded balanced domains passing through its center, namely the origin in $\mathbb{C}^{N}$. This itself is noteworthy; if not for anything else, this suffices for an important application to attain (employing the above theorem 1.1), the fairly recent result of [JJ] together with [Pranav_JJ] that: every proper holomorphic self-map of any smoothly bounded balanced pseudoconvex domain of finite type is an automorphism. We shall not however elaborate upon this here.
@@
As mentioned in the abstract, we intend a systematic presentation; so, to not be amiss about addressing the simplest of questions in the aforementioned first goal then, we recall that in [Vesntn], E. Vesentini already attained a fundamental result on the form of retracts through the center of certain convex balls. As pointed out by J. P. Vigue in [Vigue_fxdpnt_cnvx], Vesentini’s approach makes ‘heavy’ use of complex geodesics. We shall to begin with, give a relatively elementary proof of a generalization of this result without using complex geodesics. We however we emphasize that complex geodesics are conceptually significant here; they are particularly advantageous when the boundary satisfies extremality conditions, but in the presence of analytic varieties in the boundary, it seems fruitful to take a more direct approach. For instance, this approach led us among other things, to attain theorem 1.14 giving a fairly precise description of the retracts when extremality conditions are completely dropped in the setting of Vesentini’s theorem. To see what we mean by this further, we request the reader to see its proof in section 6 as well as that of proposition 1.3 below (theorem 1.17 is also based on this, though the details are longer). Perhaps more importantly, we first note that all Caratheodory geodesics (abbreviated complex $C$-geodesic) are retracts but not conversely  [^4] To mention a simple example consider the copy of $\Delta^{*}$ in $\Delta\times\Delta^{*}$, which is evidently a retract but not a complex $C$-geodesic. It may be noted in Vesentini’s theorem that the boundary is assumed to be $\mathbb{C}$-extremal which is weaker than the condition of being $\mathbb{R}$-extremal; thereby, this theorem not only applies to $\mathbb{R}$-extremal domains such as $l^{p}$-balls for $1<p<\infty$, but also more domains such as the standard $l^{1}$-ball in $\mathbb{C}^{N}$ which is not $\mathbb{R}$-extremal but indeed $\mathbb{C}$-extremal. The notions of $\mathbb{R}$-extremality, $\mathbb{C}$-extremality and the relationship between them is recalled in the following section 2 about preliminaries. The necessity of such extremality hypotheses is borne out by the fact that the assertion fails even for the polydisc and for certain simple irreducible domains as well.
@@
Here and throughout the article, by a linear subspace $D_{L}$ of a balanced domain $D\subset\mathbb{C}^{N}$, we mean that $D_{L}$ is of the form $D_{L}=D\cap L$, where $L$ is a linear subspace of $\mathbb{C}^{N}$. Likewise, by an affine linear subspace $A^{\prime}$ of $D$, we mean $A^{\prime}=A\cap D$ where $A$ is an affine linear subspace of $\mathbb{C}^{N}$. To recall a relevant fact in this language, every retract of a ball (centered at the origin, as always in this article) linearly equivalent to the standard Euclidean ball $\mathbb{B}^{N}$, is an affine linear subspace. A question which arises here is whether linear retracts $D_{L}$ can fail to be realizable as the image of a linear (retraction) map. While it is easy to ascertain that this cannot happen, it may be interesting to note that this can happen if ‘linear’ is replaced by ‘affine-linear’ i.e., there exist affine-linear retracts for certain balanced domains wherein they cannot be realized as images of affine-linear retraction maps. We postpone rendering clarity about such ambiguities and questions to section 3; likewise, for definitions of a couple of other terms as well, in order to focus more in this introductory section, on the narrative of our results and work about retracts in perspective of what is already known.
Now, generalizations of Vesentini’s theorem have certainly been obtained in the intervening years. Let us only mention a relatively recent, best result in this direction extended a bit more for our applications. In discussing this, we also attain a generalization of Mazet’s Schwarz lemma, namely theorem 3.6 of the sequel, that is a key tool for some of the main results about retracts as well as the following proposition.